# TESGNN: Temporal Equivariant Scene Graph Neural Networks for Efficient and Robust Multi-View 3D Scene Understanding

### v1

@@ Abstract
Scene graphs have proven to be highly effective for various scene understanding tasks due to their compact and explicit representation of relational information. However, current methods often overlook the critical importance of preserving symmetry when generating scene graphs from 3D point clouds, which can lead to reduced accuracy and robustness, particularly when dealing with noisy, multi-view data. This work, to the best of our knowledge, presents the first implementation of an Equivariant Scene Graph Neural Network (ESGNN) to generate semantic scene graphs from 3D point clouds, specifically for enhanced scene understanding. Furthermore, a significant limitation of prior methods is the absence of temporal modeling to capture time-dependent relationships among dynamically evolving entities within a scene. To address this gap, we introduce a novel temporal layer that leverages the symmetry-preserving properties of ESGNN to fuse scene graphs across multiple sequences into a unified global representation by an approximate graph-matching algorithm. Our combined architecture, termed the Temporal Equivariant Scene Graph Neural Network (TESGNN), not only surpasses existing state-of-the-art methods in scene estimation accuracy but also achieves faster convergence. Importantly, TESGNN is computationally efficient and straightforward to implement using existing frameworks, making it well-suited for real-time applications in robotics and computer vision. This approach paves the way for more robust and scalable solutions to complex multi-view scene understanding challenges. Our source code is publicly available at: https://github.com/HySonLab/TESGraph.
1 Introduction
@@
Holistic scene understanding plays a critical role in the advancement of robotics and computer vision systems [23, 29, 25]. The primary methodologies in this domain include 2D mapping, 3D reconstruction, and scene graph representation. Although 2D mapping remains a straightforward and widely adopted approach [37, 35], it inherently lacks a comprehensive spatial understanding of the environment. In contrast, 3D reconstruction techniques [20] provide richer spatial information but are resource intensive and particularly susceptible to noise interference [28]. Recent innovations, such as LMNav [46] and VLMap [14], leverage Large Language Models (LLMs) to generate semantic maps from 3D data, thereby enhancing scene understanding. However, these approaches often require substantial computational resources. As an alternative, scene graphs, particularly those utilizing Graph Neural Networks (GNNs), offer a more efficient and semantically rich representation compared to traditional 3D reconstruction techniques [6]. Initial methodologies mainly focused on generating scene graphs from sequences of 2D images. In more recent work [52, 53], researchers have explored scene graph generation by incorporating 3D features, such as depth data and point clouds. The latest advancements [54] propose hybrid approaches that integrate both 2D and 3D data to further refine scene graph representations. However, based on our analysis, existing approaches reveal two major limitations.
First, there is a noticeable lack of robust strategies for handling noisy, multi-view 3D data, which often results in inconsistencies within the generated scene graphs. These inconsistencies stem from the sensitivity of current methods to varying camera angles, ultimately failing to preserve symmetry. To address this, the symmetry-preserving property in Graph Neural Networks (GNNs) can be enhanced through: (1) employing invariant features for node and edge embeddings, and (2) leveraging Equivariant GNNs [26, 31]. Notably, prior works [53, 54] have yet to explore these symmetry-preserving techniques.
Second, existing methods lack effective mechanisms for constructing a coherent global scene graph from multiple point cloud sequences. For instance, [53] utilizes a heuristic averaging approach to merge several locally generated scene graphs into a single global representation. While conceptually straightforward, this method does not incorporate any optimization procedures, resulting in a global scene graph that is merely an unrefined aggregation of local graphs, lacking semantic coherence and structural consistency.
@@
To address these limitations, we propose our novel architecture Temporal Equivariant Scene Graph Neural Network (TESGNN) that contains two main components:
  1. 1.
    Equivariant Scene Graph Neural Network (ESGNN): Starting from a point cloud input, ESGNN generates node and edge embeddings to construct a scene graph. To achieve this, we design an Equivariant GNN architecture that integrates the Equivariant Graph Convolution Layer [45] with a Feature-wise Attention layer [53]. This design ensures both rotational and translational invariance, thereby improving the quality of scene representation while reducing the need for deeper layers and extensive computational resources.
  2. 2.
    Temporal Model for Merging Scene Graphs: Leveraging the local scene graphs produced by the ESGNN, this Temporal Model fuses them by approximately solving a graph matching problem via comparing the similarity of output node embeddings. The resulting unified global scene graph can subsequently be utilized for tasks such as multi-agent coordination and efficient navigation in dynamic environments.
In summary, our paper presents two key contributions:
  • To the best of our knowledge, we are the first to apply Equivariant Graph Neural Networks for semantic scene graph generation from 3D point clouds. We empirically demonstrate that incorporating symmetry-preserving techniques significantly enhances the quality of scene graph representations. Our method achieves superior accuracy and requires fewer training iterations compared to existing approaches.
  • We introduce a novel embedding-based temporal approach to merge multiple scene graphs using embedding similarity and graph matching. This is, to our knowledge, the first of its kind, and we believe it will inspire further research into temporal architectures to refine global scene graphs to achieve improved holistic multi-view scene understanding.
2 Related Work
Graph Representation Learning.
Graph representation learning has emerged as a fundamental approach for encoding structural information into low-dimensional embeddings, enabling the application of machine learning techniques to graph-structured data. The primary objective of this field is to capture the topology of graphs along with node and edge attributes [36, 11, 39, 16], facilitating tasks such as node classification, link prediction, and graph classification. Works in this domain have introduced various methods, ranging from earlier spectral techniques to recent advancements in GNNs. Graph Transformers [56, 22, 4, 32] and Graph Attention Networks [50] extend the principles of traditional transformers and attention mechanisms to graph structures, enabling the capture of long-range interactions between nodes and enhancing the model’s ability to learn complex relationships in graph data. Graph representation learning is crucial for scene graph understanding as it efficiently captures the spatial relationships among entities, providing a compact and structured representation of the environment without the need for dense point cloud data.
Equivariant Neural Networks.
While Graph Neural Networks (GNNs) are inherently permutation-equivariant [17, 27], they do not inherently preserve rotational symmetry, which is crucial for accurately capturing 3D transformations such as translations, rotations, and reflections. Achieving rotational equivariance is particularly vital when working with 3D data formats like point clouds [49], 3D meshes [2], molecular structures [1, 40], and for Neural Relational Inference in systems with interacting entities [24]. Recent advancements have focused on enforcing SE(3) or E(3) symmetries in neural networks [47, 1, 10, 9, 26]. Notably, Satorras et al. [45] introduced the E(n)-equivariant GNN (EGNN), a powerful architecture that remains equivariant under translations, rotations, reflections, and permutations for sets of input points. Building on this, Liao and Smidt [31] developed the Equiformer, which enhances Transformer GNNs by replacing standard operations with equivariant counterparts, such as tensor products. In our work, we adopt the Equivariant Graph Convolution Layer from EGNN due to its demonstrated empirical effectiveness, integrating it into our framework for improved scene graph representation.
Scence Graph Understanding.
@@ 2 Related Work
Research on scene graphs has significantly advanced tasks in vision, natural language processing, and interdisciplinary domains such as mobile robot navigation. Originally introduced as a structure to represent object instances and their relationships within a scene [21], scene graphs effectively capture rich semantics in various data modalities, including 2D/3D images [21, 3], videos [38], and point clouds [52]. Wald et al. [52] pioneered 3DSSG, the first method for generating 3D scene graphs from point clouds, accompanied by the annotated 3RScan dataset. Building on this, Wu et al. [53] introduced Scene Graph Fusion, which employs feature-wise attention to improve scene graph representations, followed by an incremental approach [54] that integrates both RGB image sequences and sparse 3D point clouds. Scene Graph Fusion remains the state-of-the-art method for scene graph generation from 3D point clouds. Our approach builds upon this by integrating Feature-wise Attention to optimize the Message-Passing process, thereby enhancing the quality of the generated scene graphs.
Temporal Graph Learning.
@@
Temporal graph learning has gained significant attention in recent years, driven by the need to model dynamically evolving relationships over time with various applications in traffic prediction [55, 30, 34] and pandemic modeling [18, 48, 33]. Traditional Graph Neural Networks (GNNs) often assume static graph structures, which limits their applicability in scenarios where graph nodes and edges change continuously, such as in video analysis and multi-view scene understanding. To address these limitations, Temporal GNNs including [43, 44] incorporate temporal dependencies and time-series learning to capture time-evolving patterns and enhance predictive performance on dynamic graphs. Our work builds upon these concepts by introducing a temporal layer that preserves symmetry, leveraging the equivariant properties of our Equivariant Scene Graph Neural Network (ESGNN). Unlike prior approaches that treat temporal information as sequential snapshots, we propose a graph matching mechanism to merge scene graphs from multiple time steps into a unified global representation, which preserves temporal consistency while reducing computational overhead.
Large-scale Point Cloud Reconstruction.
@@
3 Overall Framework
Fig. 2 illustrates the proposed framework’s capability to iteratively estimate a global 3D semantic scene graph using sequence of point clouds. The framework can be divided into two main parts, which are the feature extraction (a-c) mentioned in section 3.2 and scene graph generation (d-h) mentioned in 3.3.
Figure 2: Overview of our proposed Temporal Equivariant Scene Graph framework. Our approach takes sequences of point clouds a) as input to generate a geometric segmentations b). Subsequently, the properties of each segment and a neighbor graph between segments are constructed. The properties d) and neighbor graph e) of the segments that have been updated in the current frame c) are used as the inputs to compute node and edge features f) and to predict a 3D scene graph g). Then it goes through the temporal layer to fuse graphs from different sequences to a global one h).
3.1 Problem Formulation
Our system processes point cloud representations $D_{i}$ for each scene $i$ in a sequence $(D_{1},D_{2},\dots,D_{n})$ to generate corresponding scene graphs $(\mathcal{G}_{s,1},\mathcal{G}_{s,2},\dots,\mathcal{G}_{s,n})$ and a global graph $\mathcal{G}_{s,\text{global}}$ that aggregates scene information. Assuming effective feature extraction and geometric segmentation, our focus is on scene graph generation rather than feature extraction. Each point cloud $D_{i}$ undergoes pre-processing for input to a graph neural network (GNN). A geometric segmentation model partitions $D_{i}$ into segments, forming nodes of the scene graph, with segment attributes extracted via a point encoder. Detailed feature extraction methods are in Section 3.2.
@@ 3.2 Feature Extraction
    Geometric Segmentation and Point Cloud Extraction with Segment Nodes: Given a point cloud $D_{i}$, this geometric segmentation will provide a segment set $\mathbf{S}=\left\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\right\}$. Each segment $s_{i}$ consists of a set of 3D points $\mathbf{P}_{i}$ where each point is defined as a 3D coordinate $p_{i}\in\mathbb{R}^{3}$ and RGB color. Then, the point cloud concerning each entity is fed to the point encoder named PointNet [7] $f_{p}(\mathbf{P_{i}})$ to encode the segments $s_{i}$ into latent node and edge features, which are then passed to the model detailed in Section 4.
3.3 Scene Graph Generation
In this phase, the framework processes the input from feature extraction (Section 3.2) to generate the scene graph:
@@ 3.3 Scene Graph Generation
  • Scene Graph Representation and Temporal Graph Matching: The processed input from above is then fed to the ESGNN (Section 4) to generate the node and edge embeddings. These embeddings are used to generate the sub-graph for each sequence and merge multiple scene graphs with our Temporal Model (Section 5). For ESGNN, the node classes and edge predicates are predicted using two Multi-Layer Perceptron (MLP) classifiers. The network is trained end-to-end with a joint cross-entropy loss for both objects $\mathcal{L}_{\text{obj}}$ and predicates $\mathcal{L}_{\text{pred}}$, as described in [52]. On the other hand, our Temporal Model is trained to minimize the representation distance between identical nodes with contrastive loss from [12].
Figure 3: Overview of the ESGNN (Equivariant Scene Graph Neural Network) Architecture. The model comprises two main layers: (1) Feature-wise Attention Graph Convolution Layer (FAN-GCL) and (2) Equivariant Graph Convolution Layer (EGCL). FAN-GCL handles large inputs with multi-head attention to update node and edge features, while EGCL ensures symmetry preservation by incorporating bounding box coordinates into the message-passing mechanism. ESGNN leverages these layers to maintain rotation and permutation equivariance, and enhance the quality of scene graph generation from 3D point clouds while reducing computational costs.
4 Equivariant Scene Graph Neural Network (ESGNN)
@@ 4.1 Graph Initialization
4.2 Graph Neural Network Architecture and Weights Updating
@@
5 Temporal Model for Merging Scene Graphs
Figure 4: Temporal Matching Pipeline. The process begins with node and edge embeddings derived from different sequences (a). For each sequence, edge embeddings are concatenated with the target node embeddings to create Predicate-Object embeddings (b), which then pass through a linear layer followed by sum pooling (c). For each node, the embeddings from all associated edges are concatenated and processed through another linear layer and self-attention mechanism to generate the final embedding representation for each segment (d). These final embeddings are utilized for top K graph matching (e).
@@ 5 Temporal Model for Merging Scene Graphs
Figure 4 illustrates our ESGNN pipeline for graph matching and merging. Since sequences may differ in the number of nodes and edges, we define a Node-Edge representation for each node and its connected edges. Our goal is to produce consistent representations for identical node-edge embeddings appearing across sequences.
@@
where $(Y,X_{1},X_{2})$ are pairs of node-edge embeddings, $D_{W}$ is a distance function, and $m>0$ is a margin. We use Siamese Cosine Similarity [42] for $D_{W}$ and train with Online Contrastive Loss, focusing on hard positive and hard negative pairs. These are widely-used methods in Sentence Transformer [41], which inspire our approach for the Temporal Model.
6 Experiment
@@
Figure 5: (Comparison of recall result over training steps, column-wise interpretation. (a), (d) illustrates the result of our model versus SGFN. (b), (e) shows the ablation study. (c), (f) shows result while applying our model with the image encoder.
We evaluate TESGNN on the 3DSSG dataset [52] and compare the results with state-of-the-art works. Section 6.2 provides our model performance for scene graph generation from 3D point clouds data, in comparison to 3DSSG [52] and Scene Graph Fusion (SGFN) [53]. Our method is evaluated on full scenes given geometric segments mentioned in section 3.2. Section 6.4 discusses an extended version of our model incorporating both 2D image sequences and point clouds, built upon JointSSG [54]. Section 6.5 reports our Temporal Model for graph matching.
6.1 Dataset and Metrics
Dataset: We use the 3DSSG [52] - a dataset for scene graph generation built upon 3RScan[51], adapting the setting from [53]  [^1] https://github.com/ShunChengWu/3DSSG. The scene graph ground truths in 3DSSG are divided into 2 versions: l20, which includes 20 objects and 8 predicates, and l160, which includes 160 objects and 26 predicates. We mainly use the test set of the l20 version for our experiment as the l160 is currently having some issues within the dataset.. The original 3RScan contains 1482 3D reconstructions/snapshots of 478 naturally changing indoor environments. After processed with ScanNet [8] for geometric segmentation and grouth truth scene graph generation, the final dataset consists of 1061 sequences from 382 scenes for train, 157 sequences from 47 scenes for validation, and 117 sequences from 102 scenes for test.

### v2

@@ Abstract
Scene graphs have proven to be highly effective for various scene understanding tasks due to their compact and explicit representation of relational information. However, current methods often overlook the critical importance of preserving symmetry when generating scene graphs from 3D point clouds, which can lead to reduced accuracy and robustness, particularly when dealing with noisy, multi-view data. Furthermore, a major limitation of prior approaches is the lack of temporal modeling to capture time-dependent relationships among dynamically evolving entities in a scene. To address these challenges, we propose Temporal Equivariant Scene Graph Neural Network (TESGNN), consisting of two key components: (1) an Equivariant Scene Graph Neural Network (ESGNN), which extracts information from 3D point clouds to generate scene graph while preserving crucial symmetry properties, and (2) a Temporal Graph Matching Network, which fuses scene graphs generated by ESGNN across multiple time sequences into a unified global representation using an approximate graph-matching algorithm. Our combined architecture TESGNN outperforms current state-of-the-art methods in scene graph generation, achieving higher accuracy and faster training convergence. Moreover, we show that leveraging the symmetry-preserving property produces a more stable and accurate global scene representation compared to existing approaches. Last but not least, it is computationally efficient and easily implementable using existing frameworks, making it well-suited for real-time applications in robotics and computer vision. This approach paves the way for more robust and scalable solutions to complex multi-view scene understanding challenges. Our source code is publicly available at: https://github.com/HySonLab/TESGraph.
I Introduction
@@
Holistic scene understanding plays a critical role in the advancement of robotics and computer vision systems [1, 2, 3]. The primary methodologies in this domain include 2D mapping, 3D reconstruction, and scene graph representation. Although 2D mapping remains a straightforward and widely adopted approach [4], it inherently lacks a comprehensive spatial understanding of the environment. In contrast, 3D reconstruction techniques [5] provide richer spatial information but are resource intensive and particularly susceptible to noise interference [6]. Recent innovations, such as LMNav [7] and VLMap [8], leverage Large Language Models (LLMs) to generate semantic maps from 3D data, thereby enhancing scene understanding. However, these approaches often require substantial computational resources. As an alternative, scene graphs, particularly those utilizing Graph Neural Networks (GNNs), offer a more efficient and semantically rich representation compared to traditional 3D reconstruction techniques [9]. Initial methodologies mainly focused on generating scene graphs from sequences of 2D images. In more recent work [10, 11], researchers have explored scene graph generation by incorporating 3D features, such as depth data and point clouds. The latest advancements [12] propose hybrid approaches that integrate both 2D and 3D data to further refine scene graph representations. However, based on our analysis, existing approaches reveal two major limitations.
Current methods struggle with noisy, multi-view 3D data, leading to inconsistent scene graphs due to sensitivity to camera angles and poor symmetry preservation. Enhancing Graph Neural Networks (GNNs) with invariant node and edge embeddings and using Equivariant GNNs [13, 14] can address this, though prior works [11, 12] haven’t explored these approaches. Additionally, existing methods lack robust strategies for building coherent global scene graphs from multiple point cloud sequences. For example, [11] uses heuristic averaging to merge local graphs but skips optimization, resulting in global graphs that lack semantic coherence and structural consistency.
@@
We introduce the Temporal Equivariant Scene Graph Neural Network (TESGNN), a novel architecture that combines two key innovations:
  • Equivariant Scene Graph Neural Network (ESGNN) for Scene Graph Extraction: extracts scene graphs from 3D point clouds using an Equivariant GNN with Equivariant Graph Convolution and Feature-wise Attention layers, ensuring rotational and translational invariance. This design delivers high-quality scene representations with superior accuracy, fewer training iterations, and reduced computational demands compared to existing methods—marking the first application of symmetry-preserving Equivariant GNNs for this task.
  • Temporal Scene Graph Matching for Global Scene Representation: fuses the local scene graphs into a unified global representation by solving a graph matching problem based on node embedding similarity, with ESGNN as the backbone scene graph extractor. This embedding-based temporal approach achieves robust multi-view scene understanding, outperforming other extractors and enabling applications like multi-agent coordination and navigation in dynamic environments.
II Related Work
Graph Representation and Equivariance
Graph representation learning encodes structural and relational data into low-dimensional embeddings using methods like GNNs, Graph Transformers [15, 16, 17], and Graph Attention Networks [18], enabling tasks such as scene graph understanding without dense 3D data. For 3D applications, standard GNNs lack rotational symmetry; recent SE(3)/E(3)-equivariant architectures like EGNN [19] and Equiformer [14] address this by preserving geometric symmetries during transformations, critical for point clouds [20], molecular structures, and spatial reasoning.
Scence Graph Understanding
@@ II Related Work
Research on scene graphs has significantly advanced tasks in vision, natural language processing, and interdisciplinary domains such as mobile robot navigation. Originally introduced as a structure to represent object instances and their relationships within a scene [21], scene graphs effectively capture rich semantics in various data modalities, including 2D/3D images [21, 22], videos [23], and point clouds [10]. [10] pioneered 3DSSG, the first method for generating 3D scene graphs from point clouds, accompanied by the annotated 3RScan dataset. Building on this, [11] introduced Scene Graph Fusion, which employs feature-wise attention to improve scene graph representations, followed by an incremental approach [12] that integrates both RGB image sequences and sparse 3D point clouds. Scene Graph Fusion remains the state-of-the-art method for scene graph generation from 3D point clouds. Our approach builds upon this by integrating Feature-wise Attention to optimize the Message-Passing process, thereby enhancing the quality of the generated scene graphs.
Temporal Graph Learning
@@
Temporal graph learning has gained significant attention for modeling dynamic relationships over time, with applications in traffic prediction [24, 25, 26]. Traditional GNNs assume static graph structures, limiting their applicability in scenarios like video analysis and multi-view scene understanding. To address this, Temporal GNNs [27, 28] incorporate time-series learning to capture evolving patterns. In scene graph matching, methods such as SGAligner [29] and SG-PGM [30] align nodes across graphs but rely on ground truth graphs, making them impractical for real-world tasks where predicted graphs contain noise, ambiguous edges, and node permutations. To overcome these challenges, we introduce a symmetry-preserving temporal layer in ESGNN, leveraging equivariant properties to merge scene graphs across time steps. Unlike prior approaches that treat temporal information as sequential snapshots, our method integrates graph matching to construct a unified representation that maintains temporal consistency while reducing computational overhead.
Figure 2: Overview of our proposed TESGNN. Our approach takes sequences of point clouds a) as input to generate a geometric segmentations b). Subsequently, the properties of each segment and a neighbor graph between segments are constructed. The properties d) and neighbor graph e) of the segments that have been updated in the current frame c) are used as the inputs to compute node and edge features f) and to predict a 3D scene graph g). Then it goes through the temporal layer to fuse graphs from different sequences to a global one h).
Large-scale Point Cloud Reconstruction
@@
III Overall Framework
Fig. 2 illustrates the proposed framework’s capability to iteratively estimate a global 3D semantic scene graph from sequence of point clouds. The framework consists of three key phases: Feature Extraction (a-c), discussed in Section III-B; Scene Graph Extraction (d-f); and Temporal Graph Matching (g-h), detailed in Section III-C. Our main works, scene graph extraction and temporal graph matching, are further elaborated in Sections IV and V.
III-A Problem Formulation
Our system processes point cloud representations $D_{i}$ for each scene $i$ in a sequence $(D_{1},D_{2},\dots,D_{n})$ to generate corresponding scene graphs $(\mathcal{G}_{s,1},\mathcal{G}_{s,2},\dots,\mathcal{G}_{s,n})$ and a global graph $\mathcal{G}_{s,\text{global}}$ that aggregates scene information. Assuming effective feature extraction and geometric segmentation, our focus is on scene graph generation rather than feature extraction. Each point cloud $D_{i}$ undergoes pre-processing for input to a graph neural network (GNN). A geometric segmentation model partitions $D_{i}$ into segments, forming nodes of the scene graph, with segment attributes extracted via a point encoder. Detailed feature extraction methods are in Section III-B.
@@ III-B Feature Extraction
    Geometric Segmentation and Point Cloud Extraction with Segment Nodes: Given a point cloud $D_{i}$, this geometric segmentation will provide a segment set $\mathbf{S}=\left\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\right\}$. Each segment $s_{i}$ consists of a set of 3D points $\mathbf{P}_{i}$ where each point is defined as a 3D coordinate $p_{i}\in\mathbb{R}^{3}$ and RGB color. Then, the point cloud concerning each entity is fed to the point encoder named PointNet [36] $f_{p}(\mathbf{P_{i}})$ to encode the segments $s_{i}$ into latent node and edge features, which are then passed to the model detailed in Section IV.
III-C Scene Graph Extraction and Temporal Graph Matching
In this phase, the framework processes the input from feature extraction (Section III-B) to generate the scene graph:
@@ III-C Scene Graph Extraction and Temporal Graph Matching
  • Scene Graph Extraction and Temporal Graph Matching: The processed input from above is then fed to the ESGNN (Section IV) to generate the node and edge embeddings. These embeddings are used to generate the sub-graph for each sequence and merge multiple scene graphs with our Temporal Model (Section V). For ESGNN, the node classes and edge predicates are predicted using two Multi-Layer Perceptron (MLP) classifiers. ESGNN is trained end-to-end with a joint cross-entropy loss for both objects $\mathcal{L}_{\text{obj}}$ and predicates $\mathcal{L}_{\text{pred}}$, as described in [10]. Meanwhile, our Temporal Graph Matching is trained to minimize the representation distance between identical nodes with contrastive loss [37].
Figure 3: ESGNN Scene Graph Extractor pipeline. The model comprises two main layers: (1) Feature-wise Attention Graph Convolution Layer (FAN-GCL) and (2) Equivariant Graph Convolution Layer (EGCL). FAN-GCL handles large inputs with multi-head attention to update node and edge features, while EGCL ensures symmetry preservation by incorporating bounding box coordinates into the message-passing mechanism. ESGNN leverages these layers to maintain rotation and permutation equivariance, thus enhance the quality of scene graph generation.
IV Equivariant Scene Graph Neural Network (ESGNN)
@@ IV-A Graph Initialization
IV-B Graph Neural Network Architecture and Weights Updating
@@
V Temporal Graph Matching Network
Figure 4: Temporal Graph Matching pipeline. First, node and edge embeddings derived from scene graphs of different sequences (a). For each sequence, edge embeddings are concatenated with the target node embeddings to create Predicate-Object embeddings (b), which then pass through a linear layer followed by sum pooling (c). For each node, the embeddings from all associated edges are concatenated and processed through another linear layer and self-attention mechanism to generate the final representation of each segment (d). These final embeddings are utilized for top-K retrieval graph matching (e).
@@ V Temporal Graph Matching Network
Figure 4 illustrates our graph matching. Since sequences may differ in the number of nodes and edges, we define a Node-Edge representation for each node and its connected edges. Our goal is to produce consistent representations for identical node-edge embeddings appearing across sequences.
@@
where $(Y,X_{1},X_{2})$ are pairs of node-edge embeddings, $D_{W}$ is a distance function, and $m>0$ is a margin. We use Siamese Cosine Similarity [38] for $D_{W}$ and train with Online Contrastive Loss, focusing on hard positive and hard negative pairs. These are widely-used methods in Sentence Transformer [38], which inspire our approach.
@@
Figure 5: Comparison of recall result over training steps, column-wise interpretation. (a), (d) illustrates the result of our model versus SGFN. (b), (e) shows the ablation study. (c), (f) shows result while applying our model with the image encoder.
VI Experiment
We evaluate TESGNN on the 3DSSG dataset [10] and compare the results with state-of-the-art works. Section VI-C provides result for scene graph generation from 3D point clouds, in comparison to 3DSSG [10] and Scene Graph Fusion (SGFN) [11]. Our method is evaluated on full scenes given geometric segments mentioned in Section III-B. Section VI-E reports our Temporal Graph Matching.
VI-A Dataset
Scene Graph Extraction: We use the 3DSSG [10]  [^1] https://github.com/ShunChengWu/3DSSG. The scene graph ground truth in 3DSSG contains 2 versions: l20, which includes 20 objects and 8 predicates, and l160, which includes 160 objects and 26 predicates. We mainly use the test set of the l20 version for our experiment as the l160 is currently having some issues within the dataset, as confirmed by the authors. - a popular dataset for 3D scene graph evaluation built upon 3RScan[35], adapting the setting from SGFN [11]. The original 3RScan contains 1482 3D reconstructions/snapshots of 478 naturally changing indoor environments. After processed with ScanNet [39] for geometric segmentation and ground truth scene graph generation, the final dataset consists of 1061 sequences from 382 scenes for training, 157 sequences from 47 scenes for validation, and 117 sequences from 102 scenes for test.
Temporal Graph Matching: While reproducing prior works including SGAligner [29] and SG-PGM [30], we identified limitations in their dataset. Since their work does not contain scene graph extraction, they synthesize scene graph node and edge features using one-hot encoding of the ground truth labels, which does not accurately reflect scene graph embeddings. This set-up impractically ignores real-world noises affecting local scene graphs. To address these shortcomings, we leverage our above set-up for scene graph extraction to evaluate scene graph matching, using the same sequences and scenes. In this context, a positive pair consists of the same object appearing in two or more different sequences. There are 16,156 positive pairs over 321,292 total pairs for train set, and 3,062 positive pairs over 63,890 total pairs for test set. In our set-up, node and edge features are derived from frozen scene graphs such as ESGNN (ours) or SGFN [11].
VI-B Metrics